xx TestTypes

TestTypes

Erik Schnetter schnetter@uni-tuebingen.de

December 29 2002

Abstract

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References

6 Parameters

11 pboolean 10 Scope: private01  BOOLEAN

11 Description: boolean parameter

11          10              01  Default: yes

11 pint  10 Scope: private01  INT

11 Description: int parameter

11 Range 10              01  Default: 1

11 *:*  10

11

11 preal  10 Scope: private01  REAL

11 Description: real parameter

11 Range 10              01  Default: 1.0

11 *:*  10

11

11 pstring 10 Scope: private01  STRING

11 Description: string parameter

11 Range 10              01  Default: a

11 .*  10

11

7 Interfaces

General Implements: testtypes

Grid Variables

7.0.1 PRIVATE GROUPS

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11 vint 10 vint 01 compact 11 0
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 INT
11 vint1 10 vint1 01 compact 11 0
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 INT1
11 vint2 10 vint2 01 compact 11 0
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 INT2
11 vint4 10 vint4 01 compact 11 0
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 INT4
11 vint8 10 vint8 01 compact 11 0
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 INT8
11 vreal 10 vreal 01 compact 11 0
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL

11

11 x Group Names x 10 x Variable Names x01 Details x 11 x
11 vreal4 10 vreal4 01 compact 11 0
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL4
11 vreal8 10 vreal8 01 compact 11 0
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 REAL8
11 vcomplex 10 vcomplex 01 compact 11 0
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 COMPLEX
11 vcomplex8 10 vcomplex8 01 compact 11 0
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 COMPLEX8
11 vcomplex16 10 vcomplex16 01 compact 11 0
11                    10                    01 dimensions 11 0
11                    10                    01 distribution 11 CONSTANT
11                    10                    01 group type 11 SCALAR
11                    10                    01 timelevels 11 1
11                    10                    01 variable type 11 COMPLEX16

8 Schedule This section lists all the variables which are assigned storage by thorn CactusTest/TestTypes. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11 Always:  11 x
11 vint vint1 vint2 vint4 vint8  11 x
11 vreal vreal4 vreal8  11 x
11 vcomplex vcomplex8 vcomplex16 11 x

11 x  11 x

11

Scheduled Functions

CCTKxINITIAL x testtypesxc x test types in c x

11 x 11 Language: 11 c
11 x 11 Type:  11 function

11

CCTKxINITIAL x testtypesxfortran x test types in fortran x

11 x 11 Language: 11 fortran
11 x 11 Type:  11 function